OpenStudy (jazzyfa30):

Contagious diseases can spread very quickly. Suppose five people are ill during the first week of an epidemic and that each person who is ill spreads the disease to four people by the end of the next week. By the end of the tenth week of the epidemic, how many people have been affected by the illness?

OpenStudy (jazzyfa30):

@mathcalculus

OpenStudy (jdoe0001):

\(\bf a_n = a_1+(n-1)d\\ \text{so the 10th term/week will be}\\ \large a_{10} = a_1+(10-1)d\) d = common ratio, in this case, each quantity of ill follks will infect 4 each, so the common ratio is 5, 5*4, (5*4)*4, (5*4*4)*4 ..... is "4"

OpenStudy (jdoe0001):

hmm, I got my equation mixed up, one sec, that's the wrong one :(

OpenStudy (jdoe0001):

\(\bf a_n = a_1r^{n-1}\\ \text{so the 10th term/week will be}\\ \huge a_{10} = a_1\times r^{(10-1)}\) r = common ratio

OpenStudy (jazzyfa30):

?????

OpenStudy (jazzyfa30):

@pengembara_bumi1

OpenStudy (anonymous):

use the formula given dear

OpenStudy (jazzyfa30):

dont understand it or know how too fill it in